OpenStudy (anonymous):

I take it no one will help me out on this problem?????????????????????? 9) Find the product of (x2 - 3x + 5) with the quotient of (18x6 - 27x5 - 9x3) ÷ 9x3.

OpenStudy (experimentx):

first find the quotient.

OpenStudy (anonymous):

I have no idea how to do this kind of math

OpenStudy (experimentx):

multiplication and division (18x6 - 27x5 - 9x3) ÷ 9x3 = ??

myininaya (myininaya):

I'm having trouble understanding your wording. Do you think you can tell us what the problem says exactly?

OpenStudy (anonymous):

that is what the problem says exactly

OpenStudy (experimentx):

i also read the question ... and i am also puzzled ;////

OpenStudy (anonymous):

this is what the two problems look like (x^2 - 3x + 5) and this one (18x^6 - 27x^5 - 9x^3) ÷ 9x^3.

myininaya (myininaya):

oh do you think with is suppose to be and?

myininaya (myininaya):

Find the product of a and b means a*b

myininaya (myininaya):

Is this what your problem actually reads? \[\text{ Find the product of } x^2-3x+5 \text{ and the quotient of } 18x^6-27x^5-9x^2 \text{ and } 9x^3\]

OpenStudy (anonymous):

first divide... \[\Large (18x^6-27x^5-9x^3)\div 9x^3=\] \[\Large \frac{(18x^6-27x^5-9x^3)}{9x^3}=\] \[\Large \frac{18x^6}{9x^3}-\frac{27x^5}{9x^3}-\frac{9x^3}{9x^3}=\] \[\Large 2x^3-3x^2-1\] now multiply: \[\Large (x^2 - 3x + 5)\cdot (2x^3-3x^2-1)=\] \[\large 2x^5-3x^4-x^2-6x^4+9x^3+3x+10x^3-15x^2-5=\] \[\large 2x^5-9x^4+19x^3-16x^2+3x-5\]